# A hybrid electromechanical phase-field and deep learning framework for predicting fracture in dielectric nanocomposites

### v1

@@ 2.1 Electromechanical phase-field fracture model
This section outlines the coupled electromechanical phase-field fracture formulation under small strain assumptions. The model describes the degradation of mechanical and electrical fields due to crack nucleation and evolution in electro-active materials. The governing equations, constitutive relations, energy functional, and weak form are presented within a variational framework suitable for modeling fracture under coupled electromechanical loading conditions. For a detailed derivation and numerical implementation, the reader is referred to [13].
@@ 2.1.2 Energy-based phase-field formulation
To capture crack initiation and propagation in electro-active materials, a phase-field approach is introduced. This method regularizes the sharp crack topology $\Gamma$ using a scalar fracture variable $\mathfrak{d}$ and couples it with the mechanical and electrical fields. The total free energy of the system is formulated to incorporate the degradation of both fields as well as the energy associated with fracture.
@@
$g_{1}(\mathfrak{d})=(1-\mathfrak{d})^{2}+\eta,$ (6)
$g_{2}(\mathfrak{d},k,n)=\frac{1-\exp\left(-k\left(1-\mathfrak{d}\right)^{n}\right)}{1-\exp\left(-k\right)},$ (7)
where $0<\eta\ll 1$ is a residual stiffness parameter to prevent ill-conditioning, $k$ and $n$ are parameters determine how the degradation function behaves.
@@
where $\boldsymbol{\epsilon}$ is the permittivity tensor of the material (typically isotropic: $\boldsymbol{\epsilon}=\epsilon\mathbf{I}$, where $\epsilon$ denotes the relative permittivity and $\mathbf{I}$ is the second-order identity tensor).
@@ 2.1.4 Weak formulation
For the electrostatic equilibrium (electric potential field), from, the variation with respect to the electric potential yields:
@@
$\delta_{\mathfrak{d}}\mathcal{E}=\int_{\Omega}\left[g_{1}^{\prime}(\mathfrak{d})\psi^{\text{mech}}_{0}\delta d+G_{c}\left(\frac{\mathfrak{d}}{\ell}\delta\mathfrak{d}+\ell\nabla\mathfrak{d}\cdot\nabla\delta\mathfrak{d}\right)\right]\,\mathrm{d}\Omega.$ (31)
@@
$\boxed{\int_{\Omega}\left[g_{1}^{\prime}(\mathfrak{d})\psi^{\text{mech}}_{0}\delta\mathfrak{d}+G_{c}\left(\frac{\mathfrak{d}}{\ell}\delta\mathfrak{d}+\ell\nabla\mathfrak{d}\cdot\nabla\delta\mathfrak{d}\right)\right]\,\mathrm{d}\Omega=0\quad\forall\,\delta d\in\mathcal{W}_{d}}$ (32)
The weak forms above are discretized using the finite element method (FEM), and the resulting system of equations is implemented in the general-purpose FEM software ABAQUS using a monolithic solution strategy. For the sake of brevity, the details of FEM discretization and implementation are omitted here. For further details, the reader is referred to [13, 26].
@@ 2.2 Machine learning framework for crack segmentation
To create an efficient surrogate model capable of predicting complex fracture paths, we developed a deep learning framework based on a specialized convolutional neural network (CNN). The chosen architecture is a U-Net, which is particularly well-suited for image-to-image translation tasks such as semantic segmentation, where the output is a pixel-wise classification map of the input. This section provides a detailed description of the network’s architecture, its components, and the underlying principles that make it effective for this application.
@@ 2.2.1 Architectural overview: The U-Net model
The U-Net architecture, first proposed for biomedical image segmentation by Ronneberger et al. [43], is characterized by its symmetric encoder-decoder structure, often visualized in a ‘U’ shape shape. This design excels at capturing both high-level contextual information and fine-grained spatial details, which are essential for precisely localizing features like crack boundaries. The full architecture is illustrated in Figure 1.
The input to the network is a single-channel grayscale image of size $572\times 572$ pixels. The left side of the U is the encoder (or contracting path), which functions as a feature extractor. It consists of repeated blocks of two $3\times 3$ convolutional layers (with no padding), each followed by a rectified linear unit (ReLU) activation function. After each block, a $2\times 2$ max pooling operation with a stride of 2 is applied, halving the spatial dimensions while doubling the number of feature channels. This progressively increases the receptive field, allowing the network to build a hierarchical understanding of the image. The right side of the U is the decoder (or expansive path), which reconstructs the segmentation map from the compressed features. Each stage begins with a $2\times 2$ transposed convolution (up-convolution) that doubles the spatial resolution and halves the number of channels, followed by concatenation with the corresponding feature map from the encoder path. This is followed by two $3\times 3$ convolutional layers with ReLU activation. At the bottom of the U is the bottleneck, which marks the transition between encoding and decoding. It consists of a convolutional block that processes the most abstracted representation of the input. The final layer applies a $1\times 1$ convolution to reduce the number of output channels to two, producing a $388\times 388$ segmentation map. A softmax activation is applied to obtain pixel-wise classification probabilities.
Figure 1: Schematic overview of the U-Net architecture employed for crack segmentation, adapted from [43].
@@ 2.2.2 Feature extraction enhancement via ResNet backbone
Second, the ResNet architecture itself is fundamentally more powerful for training very deep networks. Its core innovation is the ”residual block,” where the output of a layer (or a stack of layers) is added to its original input via a ”skip” or ”identity” connection. This allows the network to learn residual functions, effectively letting gradients flow more easily through the deep network during backpropagation. This structure elegantly solves the vanishing gradient problem that plagues traditional deep CNNs, enabling the construction of much deeper and more expressive models. We systematically evaluated four ResNet configurations–ResNet34, ResNet50, ResNet101, and ResNet152–as the encoder to analyze the trade-off between model complexity, training time, and predictive accuracy.
@@ 2.3 Dataset generation
A comprehensive and physically grounded dataset is developed to train and validate deep learning models for crack path prediction in smart electromechanical materials. The data generation process is rooted in high-fidelity simulations based on the fully coupled electromechanical phase-field fracture model presented above. The resulting dataset encapsulates the evolution of displacement, electric potential, and fracture damage fields across a range of mechanical and electrical loading scenarios. This section details the simulation framework, data structure, and augmentation strategies employed to generate a diverse and generalizable image-based dataset.
@@ 2.3.1 Simulation framework
The dataset was generated using ABAQUS in combination with a custom-developed user-defined element (UEL) subroutine that implements a fully coupled electromechanical phase-field fracture formulation. This framework simultaneously solves for the displacement field, electric potential, and phase-field fracture variable, allowing accurate modeling of electromechanically induced fracture in heterogeneous media.
@@
Table 1:
Material properties of the boehmite nanoparticle-reinforced polymer composite used in simulations, see [44, 45]. Property Symbol Value Young’s modulus $E$ 3.75 GPa Poisson’s ratio $\nu$ 0.33 Relative permittivity $\epsilon$ 5.5 Fracture toughness $\mathcal{G}_{c}$ 165 J/m2 Residual stiffness parameter $\eta$ $1\times 10^{-4}$ Degradation threshold $k$ $1\times 10^{-4}$ Degradation exponent $n$ 2.0
The simulated specimen consists of a square plate measuring 1 mm $\times$ 1 mm with a thickness of 0.05 mm, representative of a thin-film nanocomposite material. To introduce geometric variability and simulate defect-induced failure mechanisms, each model incorporates 10 to 15 randomly distributed circular holes. The hole diameters range from 0.02 mm to 0.06 mm, with a minimum inter-hole spacing of 0.03 mm to avoid overlap and preserve structural integrity. A mesh convergence study was conducted to ensure the quality and accuracy of the finite element discretization prior to running the full simulation set. Hole placement and sizing were automated using a custom Python script integrated within ABAQUS, enabling efficient generation of spatially randomized but reproducible geometries across 10,000 simulation instances.
@@
In terms of loading, a displacement-controlled test was simulated by prescribing a total displacement of 0.15 mm at the right edge, while fixing the left edge in the horizontal direction. For the electromechanical coupling, Dirichlet boundary conditions were applied by prescribing a potential difference across the specimen: 0 V at the left edge and 1 V at the right edge. To prevent numerical artifacts at the boundaries, all holes were positioned at least 0.1 mm away from any plate edge.
@@ 2.3.2 Dataset composition
Figure 3: Representative simulations – (Left) Reference/original configuration, (Middle) Phase-field crack path, and (Right) Electrical potential crack path.
@@ 3.1 Evaluation metrics
To quantitatively assess the predictive accuracy of the CNN models across different input modalities–namely, phase-field fracture and electrical potential images–a set of core evaluation metrics was employed. The F1-score was used to balance precision and recall, making it particularly effective for binary segmentation tasks where there is a class imbalance between fractured and unfractured regions. Intersection over Union (IoU) was adopted to quantify the spatial overlap between the predicted crack regions and the ground truth, providing a robust measure of localization performance. To evaluate model convergence, a combination of binary cross-entropy and dice loss functions was used, with lower values indicating closer alignment between the predicted and actual segmentation maps. Additionally, validation metrics–including validation F1-score (Val F1), validation IoU (Val IoU), and validation loss (Val Loss)–were tracked to assess the models’ generalization capabilities on unseen data. Finally, training time was recorded to evaluate computational efficiency, which is critical for practical deployment in real-time or resource-constrained environments. All metrics were computed per epoch and averaged across the training and validation datasets to ensure statistical consistency and robust model assessment.

### v2

@@ 2.1 Electromechanical phase-field fracture model
This section presents a coupled electromechanical phase-field fracture formulation under small-strain assumptions, describing the degradation of mechanical and electrical fields due to crack nucleation and evolution in electro-active materials. The governing equations, constitutive relations, energy functional, and weak form are formulated within a variational framework for coupled electromechanical fracture. Detailed derivations and numerical implementation are referred to Quinteros et al. [2023].
@@ 2.1.2 Energy-based phase-field formulation
Crack initiation and propagation in electro-active materials are modeled using a phase-field approach, in which the sharp crack topology $\Gamma$ is regularized by the scalar fracture variable $\mathfrak{d}$ and coupled to the mechanical and electrical fields. The total free energy accounts for the degradation of both fields as well as the fracture energy.
@@
$g_{1}(\mathfrak{d})=(1-\mathfrak{d})^{2}+\kappa_{\eta},$ (6)
$g_{2}(\mathfrak{d},k,n)=\frac{1-\exp\left(-k\left(1-\mathfrak{d}\right)^{n}\right)}{1-\exp\left(-k\right)}+\kappa_{\epsilon},$ (7)
where $\kappa_{\eta}$ and $\kappa_{\epsilon}$ are residual parameters to prevent ill-conditioning, $k$ and $n$ are parameters that determine how the degradation function behaves. The exponential dielectric degradation function $g_{2}(\mathfrak{d})$ is adopted from Quinteros et al. [2023], where it was shown to effectively capture the nonlinear reduction of permittivity observed in electroactive and dielectric solids under damage. Unlike commonly used quadratic forms, this two-parameter formulation provides additional flexibility through $k$ and $n$ to control the onset and rate of dielectric degradation independently of the mechanical response. The function satisfies the essential physical requirements, ensuring boundedness and a smooth transition between intact and fully fractured states.
The crack in this framework is represented by the phase-field variable $\mathfrak{d}\in[0,1]$, which degrades both the elastic stiffness and the dielectric permittivity through the degradation function $\boldsymbol{\epsilon}(\mathfrak{d})=g_{2}(\mathfrak{d})\boldsymbol{\epsilon}_{0}$, where $g_{2}(1)=\kappa_{\epsilon}\ll 1$. Because the degraded permittivity satisfies $\kappa_{\epsilon}>0$, the electric field inside fully damaged regions ($\mathfrak{d}\approx 1$) does not vanish entirely; instead, a small residual electric field may penetrate the crack volume. Physically, this corresponds to a crack that is not a perfect vacuum gap but rather a partially bridged or micro-cracked zone, which is common in nanocomposites due to filler ligaments, see Tan et al. [2023], Abdollahi and Arias [2012].
@@
where $\boldsymbol{\epsilon}$ is the permittivity tensor of the material (typically isotropic: $\boldsymbol{\epsilon}=\epsilon\mathbf{I}$, where $\epsilon$ denotes the dielectric permittivity and $\mathbf{I}$ is the second-order identity tensor).
@@ 2.1.3 Variational principles and Euler–Lagrange equations
The interactions between the mechanical, electrical, and damage fields are summarized in Table 1. Physically, the elastic strain energy density acts as the primary driving force for the evolution of the phase-field. As damage accumulates, the local degradation of the material stiffness and dielectric permittivity redistributes the mechanical stresses and alters the electric potential distribution.
Table 1: Summary of field variables, governing equations, and coupling mechanisms in the electromechanical phase-field framework.
Physical field Variable Governing equation Boundary conditions
Mechanical $\mathbf{u}$ Momentum balance (Eq. 12) Prescribed $\bar{\mathbf{u}}$ or $\bar{\mathbf{t}}$
Electrical $\phi$ Gauss’s law (Eq. 13) Prescribed $\bar{\phi}$ or $\bar{q}_{e}$
Damage $\mathfrak{d}$ Phase-field evolution (Eq. 14) No-flux ($\nabla\mathfrak{d}\cdot\mathbf{n}=0$)
Coupling Physical Mechanism
Mech. $\to$ Damage Elastic energy $\psi_{0}^{\text{mech}}$ acts as the driving force for $\mathfrak{d}$ evolution.
Damage $\to$ Mech. Stiffness tensor $\mathbb{C}$ is degraded by $g_{1}(\mathfrak{d})$, altering the stress state.
Damage $\to$ Elec. Permittivity $\boldsymbol{\epsilon}$ is degraded by $g_{2}(\mathfrak{d})$, imposing insulating conditions.
@@ 2.1.4 Weak formulation
For the electrostatic equilibrium (electric potential field), the variation with respect to the electric potential yields:
@@
$\delta_{\mathfrak{d}}\mathcal{E}=\int_{\Omega}\left[g_{1}^{\prime}(\mathfrak{d})\psi^{\text{mech}}_{0}\delta d{\color[rgb]{0,0,0}+g_{2}^{\prime}(\mathfrak{d})\psi^{\text{elec}}_{0}\delta\mathfrak{d}}+G_{c}\left(\frac{\mathfrak{d}}{\ell}\delta\mathfrak{d}+\ell\nabla\mathfrak{d}\cdot\nabla\delta\mathfrak{d}\right)\right]\,\mathrm{d}\Omega,$ (31)
where $\psi_{0}^{\text{elec}}=\frac{1}{2}\mathbf{E}\cdot\boldsymbol{\epsilon}\cdot\mathbf{E}$.
@@
$\boxed{\int_{\Omega}\left[g_{1}^{\prime}(\mathfrak{d})\psi^{\text{mech}}_{0}\delta\mathfrak{d}{\color[rgb]{0,0,0}+g_{2}^{\prime}(\mathfrak{d})\psi^{\text{elec}}_{0}\delta\mathfrak{d}}+G_{c}\left(\frac{\mathfrak{d}}{\ell}\delta\mathfrak{d}+\ell\nabla\mathfrak{d}\cdot\nabla\delta\mathfrak{d}\right)\right]\,\mathrm{d}\Omega=0\quad\forall\,\delta d\in\mathcal{W}_{d}}$ (32)
The weak forms above are discretized using the finite element method (FEM), and the resulting system of equations is implemented in the general-purpose FEM software ABAQUS using a monolithic solution strategy. For the sake of brevity, the details of FEM discretization and implementation are omitted here. For further details, the reader is referred to Quinteros et al. [2023], Asur Vijaya Kumar et al. [2022].
@@ 2.2 Machine learning framework for crack segmentation
To create an efficient surrogate model capable of predicting complex fracture paths, we developed a deep learning framework based on a specialized convolutional neural network. The chosen architecture is a U-Net, which is particularly well-suited for image-to-image translation tasks such as semantic segmentation, where the output is a pixel-wise classification map of the input. This section provides a detailed description of the network’s architecture, its components, and the underlying principles that make it effective for this application.
@@ 2.2.1 Architectural overview: The U-Net model
The U-Net architecture, first proposed for biomedical image segmentation by Ronneberger et al. Ronneberger et al. [2015], is characterized by its symmetric encoder-decoder structure, often visualized in a ‘U’ shape. This design excels at capturing both high-level contextual information and fine-grained spatial details, which are essential for precisely localizing features like crack boundaries. The full architecture is illustrated in Figure 1.
The input to the network is a single-channel grayscale image of size $572\times 572$ pixels. The left side of the U is the encoder (or contracting path), which functions as a feature extractor. It consists of repeated blocks of two $3\times 3$ convolutional layers (with no padding), each followed by a rectified linear unit (ReLU) activation function. After each block, a $2\times 2$ max pooling operation with a stride of 2 is applied, halving the spatial dimensions while doubling the number of feature channels. This progressively increases the receptive field, allowing the network to build a hierarchical understanding of the image.
The right side of the U is the decoder (or expansive path), which reconstructs the segmentation map from the compressed features. Each stage begins with a $2\times 2$ transposed convolution (up-convolution) that doubles the spatial resolution and halves the number of channels, followed by concatenation with the corresponding feature map from the encoder path. This is followed by two $3\times 3$ convolutional layers with ReLU activation.
At the bottom of the U is the bottleneck, which marks the transition between encoding and decoding. It consists of a convolutional block that processes the most abstracted representation of the input. The final layer applies a $1\times 1$ convolution to reduce the number of output channels to two, producing a $388\times 388$ segmentation map. A softmax activation is applied to obtain pixel-wise classification probabilities.
Figure 1: Schematic overview of the U-Net architecture employed for crack segmentation, adapted from Ronneberger et al. [2015].
@@ 2.2.2 Feature extraction enhancement via ResNet backbone
Second, the ResNet architecture itself is fundamentally more powerful for training very deep networks. Its core innovation is the ”residual block,” where the output of a layer (or a stack of layers) is added to its original input via a ”skip” or ”identity” connection. This allows the network to learn residual functions, effectively letting gradients flow more easily through the deep network during backpropagation. This structure elegantly solves the vanishing gradient problem that plagues traditional deep CNNs, enabling the construction of much deeper and more expressive models. We systematically evaluated four ResNet configurations: ResNet34, ResNet50, ResNet101, and ResNet152. In this notation, the numbers (e.g., 34, 152) correspond to the total depth of the network (i.e., the number of convolutional and fully connected layers), see Ronneberger et al. [2015]. These specific architectures were selected to span a representative range of model capacities, allowing for a systematic investigation of how increasing backbone depth affects crack-segmentation performance and overfitting tendencies.
@@ 2.3 Dataset generation
All high-fidelity simulations are carried out in ABAQUS using a custom user-defined element (UEL) implementing the coupled electromechanical phase-field formulation. The nodal field outputs of the phase-field variable $\mathfrak{d}$ and electric potential $\phi$ are exported and post-processed into structured grayscale images on a uniform grid. These images form a physically grounded dataset encapsulating geometrical configurations, electric potential, and fracture damage under prescribed mechanical and electrical loading scenarios. The dataset serves as the interface between the physics-based model and the data-driven stage.
The deep learning workflow is implemented in a standalone Python environment using TensorFlow/Keras. The geometric configuration of the plate–specifically the binary distribution of holes–is used as the primary input for the CNN models. The exported physical field images from the FEM simulations (phase-field and electric potential) serve as the ground-truth targets to guide the network during the training and evaluation phases. No direct coupling with ABAQUS is required during prediction; surrogate modeling is performed offline using the pre-generated dataset. In the following sections, the simulation framework, data structure, and augmentation strategies used to generate this diverse, geometry-to-fracture mapping dataset are detailed.
@@ 2.3.1 Simulation framework
The dataset was generated using ABAQUS in combination with a custom-developed UEL subroutine that implements a fully coupled electromechanical phase-field fracture formulation. This framework simultaneously solves for the displacement field, electric potential, and phase-field fracture variable, allowing accurate modeling of electromechanically induced fracture in heterogeneous media.
@@
Table 2: Material properties of the boehmite nanoparticle-reinforced polymer composite used in simulations, see Jux et al. [2018], Fankhänel [2020].
Property Symbol Value
Young’s modulus $E$ 3.75 GPa
Poisson’s ratio $\nu$ 0.33
Dielectric permittivity $\epsilon$ $4.9\times 10^{-11}$
Fracture toughness $\mathcal{G}_{c}$ 165 J/m2
Degradation threshold $k$ $1\times 10^{-4}$
Degradation exponent $n$ 2.0
The simulated specimen consists of a square plate measuring 1 mm×1 $\text{ mm}\times 1\text${ mm} with a thickness of $0.05\text{ mm}$, representative of a thin-film nanocomposite material. To introduce geometric variability and simulate defect-induced failure mechanisms, each model incorporates 10 to 15 randomly distributed circular holes with diameters ranging from $0.02\text{ mm}$ to $0.06\text{ mm}$. This stochastic geometric setup serves as a geometry-agnostic benchmark designed to generate a statistically rich and diverse dataset of crack-defect interactions. By varying the spatial configuration of these holes, we ensure the deep learning framework learns the underlying physics of electromechanical field perturbations rather than memorizing a specific structural shape, thereby enhancing its generalizability.
A mesh convergence study was conducted to ensure the quality and accuracy of the finite element discretization prior to running the full simulation set. Hole placement and sizing were automated using a custom Python script integrated within ABAQUS, enabling efficient generation of spatially randomized but reproducible geometries across 10,000 simulation instances.
@@
In terms of loading, a displacement-controlled test was simulated by prescribing a total displacement of 0.15 mm at the right edge, while fixing the left edge. For the electromechanical coupling, Dirichlet boundary conditions were applied with a potential difference of 1 V across the specimen. These boundary conditions were simplified to isolate the fundamental electromechanical coupling and prevent the dataset from being biased toward a specific structural geometry.
This configuration is highly relevant to Structural Health Monitoring (SHM) applications. In practical applications, the electric potential represents a physically measurable quantity via electrodes; our model demonstrates that these measurable field perturbations act as a unique ’signature’ that can be mapped directly to internal fracture states. To prevent numerical artifacts, all holes were positioned at least 0.1 mm from the edges, ensuring crack initiation is driven solely by internal defect interactions.
@@ 2.3.2 Dataset composition
Figure 3: Representative simulations – (Left) Reference/original configuration, (Middle) Phase-field crack path, and (Right) Electric potential crack path. Note: These plots are shown as standard false-color (RGB) contours for illustrative clarity; the actual data utilized for the deep learning framework are single-channel grayscale images derived from these distributions.
@@
To provide a clear overview of the framework’s implementation, the process is structured into distinct offline and online stages. This separation ensures that the computational complexity of the fully coupled electromechanical simulations is confined to the data-generation phase, while the predictive phase remains computationally lightweight. Algorithm 1 outlines the step-by-step workflow, from the initial physics-based modeling to the final surrogate-based inference. Crucially, while the training process is informed by coupled physical signatures–specifically the phase-field and electric potential data–the online deployment requires only the geometric configuration as input, enabling near-instantaneous fracture assessment without the need for live field calculations.
1:Phase I: Offline high-fidelity data generation
2:Geometric initialization: Define NN samples of the domain with varying defect (hole) configurations.
3:Coupled PF-simulation: Solve the governing electromechanical equations for each sample to obtain the final fracture state.
4:Data extraction: Export the initial geometric configuration ($X$), along with the resulting phase-field ($\phi$) and electric potential field, highlighting the final crack path.
5:Phase II: Training and physics-based learning
6:Data normalization: Scale physical fields $\phi$ and $\mathfrak{d}$ to the range $[0,1]$ for numerical stability.
7:CNN training: Train the ResNet-U-Net to map geometric features to the final fracture path.
8:Feature supervision: Compare training strategies supervised by phase-field targets vs. those informed by global electric potential gradients.
9:Phase III: Predictive inference (Deployment)
10:Input reception: Input a new geometric configuration image into the trained network.
11:Surrogate execution: Perform a forward pass through the CNN (bypassing the iterative FEM solver).
12:Result output: Generate the predicted crack path mask for instantaneous assessment.
Algorithm 1 Implementation of the hybrid electromechanical DL framework.
@@ 3.1 Evaluation metrics
To quantitatively assess the predictive accuracy of the CNN models across different modalities–namely, phase-field fracture and electric potential images–a set of core evaluation metrics was employed. The F1-score was used to balance precision and recall, making it particularly effective for binary segmentation tasks where there is a class imbalance between fractured and unfractured regions. Intersection over Union (IoU) was adopted to quantify the spatial overlap between the predicted crack regions and the ground truth, providing a robust measure of localization performance. To evaluate model convergence, a combination of binary cross-entropy and dice loss functions was used, with lower values indicating closer alignment between the predicted and actual segmentation maps. Additionally, validation metrics–including validation F1-score (Val F1), validation IoU (Val IoU), and validation loss (Val Loss)–were tracked to assess the models’ generalization capabilities on unseen data. Finally, training time was recorded to evaluate computational efficiency, which is critical for practical deployment in real-time or resource-constrained environments. All metrics were computed per epoch and averaged across the training and validation datasets to ensure statistical consistency and robust model assessment.